Observation Start Time:
Observation End Time:
This night, we tested another interface to couple the spectroscope
with the telescope with a hole in the interace for the visualisation
of the fiber. Once again, the system did'nt work but the issue is now
found. I'm not sure that the issue is solvable. I hope to find a
solution before the last night
The problems was explained in the summary.